1. A 360 W appliance is connected to a 120 V power supply. What current flows through this appliance? What is the resistance of this appliance? 2. Calculate the resistance of 4.00 km of number 12 gauge copper wire at 20.c. R1 3. For the circuit below calculate: 3.0 n (a) The total resistance RT Vt= 24 V (b) The total current through the power supply. 153 15 (c) The current through each resistor. 15s (d) The voltage drop across R R2 4.0 n 4. For the circuit below, determine the currents and their direction. 6.0 4.0 51 4.0 L. 6.0 v 12.0 v I need all the work as well
Answer #1

1] we know that the power P = i*V

360 = i*120

i = 360/120 = 3 A

Resistance R = V/i = 120/3 = 40 ohm

Hence the current that flows is 3A and resistance is 40 ohm